OpenStudy (peterg2001):

Two numbers are in the ratio of 2 to 3. If the smaller number is 18, the larger number is _____. 21 27 36 54

OpenStudy (jhannybean):

\[\frac{2}{3} = \frac{18}{x}\]

OpenStudy (jhannybean):

I wrote the ratio out as \[\frac{small~ number}{big~number}\]Therefore, if the small number is ontop, we're missing the big number, \(x\).

OpenStudy (peterg2001):

27?

OpenStudy (jhannybean):

Good job :)